# Structured Self-Consistency: A Multi-Task Evaluation of LLMs on VirtualHome

### v1

@@ 1  Introduction
The VirtualHome environment (Puig et al., 2018) provides a realistic household simulation platform for evaluating embodied agents through long-horizon, goal-directed activities. When combined with the Embodied Agent Interface (EAI) (EAI, 2024), it enables a standardized and modular evaluation framework that captures multiple dimensions of embodied intelligence, as illustrated in Figure 1.
@@
Recent advances in Large Language Models (LLMs) have demonstrated strong capabilities in reasoning and task planning (Brown et al., 2020; Ahn et al., 2022). However, applying LLMs to embodied AI introduces a set of challenges that fundamentally distinguish embodied decision-making from conventional natural language processing tasks:
@@
In this work, we present a systematic evaluation of two representative LLMs, OpenPangu-7B and Qwen2.5-7B-Instruct, on the VirtualHome benchmark under the EAI framework. To address the challenges of structured generation in embodied settings, we introduce Structured Self-Consistency (SSC), an inference-time decoding framework that extends the self-consistency paradigm (Wang et al., 2023) with task-specific validation and structure-aware voting mechanisms.
@@ 2  Related Work
Embodied AI and Simulation Environments. Realistic simulation environments are pivotal for embodied AI. VirtualHome (Puig et al., 2018) supports complex household activities via programs, while AI2-THOR (Kolve et al., 2017) offers physics-enabled 3D environments for navigation and manipulation. Habitat (Savva et al., 2019) emphasizes efficiency for large-scale training. Recently, the Embodied Agent Interface (EAI) (EAI, 2024) established a standardized framework unifying task definitions across simulators. Unlike prior work focused on single tasks, we leverage EAI’s multi-task protocol to comprehensively assess LLM capabilities across the embodied intelligence spectrum.
LLMs for Robotic Task Planning. LLMs demonstrate strong potential as robotic planners. SayCan (Ahn et al., 2022) introduced affordance-aware planning, Code as Policies (Liang et al., 2023) enabled control code generation, ProgPrompt (Singh et al., 2023) produced executable programs, and Inner Monologue (Huang et al., 2022) incorporated interactive feedback. While most studies (Ahn et al., 2022; Liang et al., 2023; Singh et al., 2023; Huang et al., 2022) focus on proprietary models or end-to-end execution metrics. In contrast, we systematically compare open-source LLMs across diverse embodied tasks to reveal complementary strengths.
Self-Consistency and Decoding Strategies. Self-Consistency (Wang et al., 2023) improved reasoning via sampling and majority voting. Extensions include Universal Self-Consistency (Chen et al., 2023) for free-form outputs, Self-Verification (Weng et al., 2023) for checking, and execution-based selection (Chen et al., 2021) for code correctness. Building on these, our Structured Self-Consistency (SSC) introduces domain-specific validation and skeleton-based voting to address the unique challenges of structured generation in embodied AI.
@@ 4.1  Experimental Settings
  • Qwen2.5-7B-Instruct is a general-purpose instruction-following model trained on large-scale multilingual and code-rich corpora. Owing to its strong instruction alignment, it demonstrates robust compliance with structured output formats and stable low-level action generation. As a result, Qwen2.5-7B-Instruct typically achieves strong baseline performance on tasks requiring strict schema adherence and executable action sequences, making it a competitive generalist baseline for embodied decision-making.
  • OpenPangu-7B is a general-purpose foundation model with comparatively weaker instruction-following and formatting robustness in zero-shot settings. Its raw outputs often suffer from schema violations or surface-level inconsistencies. Nevertheless, OpenPangu-7B exhibits strong latent semantic understanding and hierarchical planning capabilities, particularly in tasks involving long-horizon reasoning and subgoal structure. This makes it an ideal testbed for examining whether structured inference frameworks can unlock the underlying reasoning potential of less instruction-aligned models.

### v2

@@ 1  Introduction
The VirtualHome environment (Puig et al., 2018) provides a realistic household simulation platform for evaluating embodied agents through long-horizon, goal-directed activities. When combined with the Embodied Agent Interface (EAI) (Li et al., 2024), it enables a standardized and modular evaluation framework that captures multiple dimensions of embodied intelligence, as illustrated in Figure 1.
@@
Recent advances in Large Language Models (LLMs) have demonstrated strong capabilities in reasoning and task planning (Mann et al., 2020; Ahn et al., 2022). However, applying LLMs to embodied AI introduces a set of challenges that fundamentally distinguish embodied decision-making from conventional natural language processing tasks:
@@
In this work, we present a systematic evaluation of two representative LLMs, OpenPangu-7B(Chen et al., 2025) and Qwen2.5-7B-Instruct(Qwen et al., 2025), on the VirtualHome benchmark under the EAI framework. To address the challenges of structured generation in embodied settings, we introduce Structured Self-Consistency (SSC), an inference-time decoding framework that extends the self-consistency paradigm (Wang et al., 2022) with task-specific validation and structure-aware voting mechanisms.
@@ 2  Related Work
Embodied AI and Simulation Environments. Realistic simulation environments are pivotal for embodied AI. VirtualHome (Puig et al., 2018) supports complex household activities via programs, while AI2-THOR (Kolve et al., 2017) offers physics-enabled 3D environments for navigation and manipulation. Habitat (Savva et al., 2019) emphasizes efficiency for large-scale training. Recently, the Embodied Agent Interface (EAI) (Li et al., 2024) established a standardized framework unifying task definitions across simulators. Unlike prior work focused on single tasks, we leverage EAI’s multi-task protocol to comprehensively assess LLM capabilities across the embodied intelligence spectrum.
LLMs for Robotic Task Planning. LLMs demonstrate strong potential as robotic planners. SayCan (Ahn et al., 2022) introduced affordance-aware planning, Code as Policies (Liang et al., 2022) enabled control code generation, ProgPrompt (Singh et al., 2022) produced executable programs, and Inner Monologue (Huang et al., 2022) incorporated interactive feedback. While most studies (Ahn et al., 2022; Liang et al., 2022; Singh et al., 2022; Huang et al., 2022) focus on proprietary models or end-to-end execution metrics. In contrast, we systematically compare open-source LLMs across diverse embodied tasks to reveal complementary strengths.
Self-Consistency and Decoding Strategies. Self-Consistency (Wang et al., 2022) improved reasoning via sampling and majority voting. Extensions include Universal Self-Consistency (Chen et al., 2023) for free-form outputs, Self-Verification (Weng et al., 2023) for checking, and execution-based selection (Chen, 2021) for code correctness. Building on these, our Structured Self-Consistency (SSC) introduces domain-specific validation and skeleton-based voting to address the unique challenges of structured generation in embodied AI.
@@ 4.1  Experimental Settings
  • Qwen2.5-7B-Instruct(Qwen et al., 2025) is a general-purpose instruction-following model trained on large-scale multilingual and code-rich corpora. Owing to its strong instruction alignment, it demonstrates robust compliance with structured output formats and stable low-level action generation. As a result, Qwen2.5-7B-Instruct typically achieves strong baseline performance on tasks requiring strict schema adherence and executable action sequences, making it a competitive generalist baseline for embodied decision-making.
  • OpenPangu-7B(Chen et al., 2025) is a general-purpose foundation model with comparatively weaker instruction-following and formatting robustness in zero-shot settings. Its raw outputs often suffer from schema violations or surface-level inconsistencies. Nevertheless, OpenPangu-7B exhibits strong latent semantic understanding and hierarchical planning capabilities, particularly in tasks involving long-horizon reasoning and subgoal structure. This makes it an ideal testbed for examining whether structured inference frameworks can unlock the underlying reasoning potential of less instruction-aligned models.